# Vibe Researching as Wolf Coming: Can AI Agents with Skills Replace or Augment Social Scientists?

### v1

@@ Abstract
AI agents—systems that execute multi-step reasoning workflows with persistent state, tool access, and specialist skills—represent a qualitative shift from prior automation technologies in social science. Unlike chatbots that respond to isolated queries, AI agents can now read files, run code, query databases, search the web, and invoke domain-specific skills to execute entire research pipelines autonomously. This paper introduces the concept of vibe researching—the AI-era parallel to “vibe coding” (Karpathy, 2025)—and uses scholar-skill, a 21-skill plugin for Claude Code covering the full research pipeline from idea to submission, as an illustrative case. I develop a cognitive task framework that classifies research activities along two dimensions—codifiability and tacit knowledge requirement—to identify a delegation boundary that is cognitive, not sequential: it cuts through every stage of the research pipeline, not between stages. I argue that AI agents excel at speed, coverage, and methodological scaffolding but struggle with theoretical originality and tacit field knowledge. The paper concludes with an analysis of three implications for the profession—augmentation with fragile conditions, stratification risk, and a pedagogical crisis—and proposes five principles for responsible vibe researching.
@@ 1 Introduction
I address this question by developing a cognitive task framework and applying it to a concrete case: scholar-skill, a 21-skill plugin for Claude Code that covers the entire social science research workflow, from initial puzzle to journal submission. The system targets publication in top journals including the American Sociological Review, the American Journal of Sociology, Demography, Science Advances, Nature Human Behaviour, and Nature Computational Science.
@@ 2.2 From Chatbot to AI Agent
An AI agent, by contrast, is multi-step, persistent, and tool-using (Park et al., 2023; Wu et al., 2023). In the context of social science research, this means the agent can: read and write files to disk; execute R, Python, or Stata code via a shell; query databases (Zotero, NHANES, IPUMS); search the web and fetch papers; spawn parallel sub-agents for independent tasks; and invoke specialist skills with domain-specific knowledge bases. This is a qualitatively different architecture that enables a qualitatively different range of tasks (Shinn et al., 2024).
@@ 3.1 Architecture and Scope
Scholar-skill is a plugin for Claude Code—Anthropic’s command-line interface for the Claude language model—that organizes 21 specialist AI skills into a complete research pipeline. The system is installed locally, runs on the researcher’s laptop, and accesses local files, databases, and code environments. Table 2 summarizes the seven stage-groups and their constituent skills.
Table 2: Scholar-Skill: 21 Skills Organized by Research Stage
@@
Each skill is a self-contained workflow with its own knowledge base: reference documents, code templates, journal-specific norms, and quality checklists. The orchestrator—scholar-full-paper—coordinates all 21 skills across 14 phases with 18 quality gates that prevent the pipeline from advancing without meeting minimum standards.
@@ 3.2 Key Technical Capabilities
Asset-driven writing. Scholar-write uses a three-tier knowledge graph built from 127 published papers: 32 by the author, 8 collaborator articles, and 87 top-journal exemplars. Tier 1 provides a section-snippets library organized by nine rhetorical functions. Tier 2 provides pre-extracted stylistic annotations (opening hooks, gap sentences, contribution claims, voice registers, citation density norms) for all 127 papers. Tier 3 provides full PDF access for deep reads. The system generates original prose that mimics the rhetorical structure of published papers while producing novel content—a distinction that separates structure mimicry from plagiarism.
Peer review simulation. Scholar-respond spawns three reviewer agents in parallel—a methods specialist, a theory expert, and a senior editor—each calibrated to the target journal’s priorities. The simulation produces a triage dashboard classifying concerns as MAJOR-FEASIBLE, MINOR-EASY, DISAGREE, or CONFLICT, with cross-reviewer overlaps flagged for priority attention.
@@ 6.2 The Stratification Risk
Cost. Claude Pro is twenty dollars a month. API tokens for a full 14-phase pipeline run cost five to fifteen dollars. For researchers at resource-poor institutions, these are meaningful barriers.
@@ 8 Discussion and Conclusion
First, AI agents are qualitatively different from prior tools. They execute multi-step reasoning workflows—not just computation or text—across the entire research pipeline. The scholar-skill system demonstrates that a single plugin can now cover 21 distinct research tasks from idea formalization to journal submission, coordinated by an orchestrator with quality gates. The change in what a solo researcher can accomplish is real and significant.

### v2

# arXiv:2602.22401v2[cs.AI] 27 Feb 2026

@@ Abstract
AI agents—systems that execute multi-step reasoning workflows with persistent state, tool access, and specialist skills—represent a qualitative shift from prior automation technologies in social science. Unlike chatbots that respond to isolated queries, AI agents can now read files, run code, query databases, search the web, and invoke domain-specific skills to execute entire research pipelines autonomously. This paper introduces the concept of vibe researching—the AI-era parallel to “vibe coding” (Karpathy, 2025)—and uses scholar-skill, a 23-skill plugin for Claude Code covering the full research pipeline from idea to submission, as an illustrative case. I develop a cognitive task framework that classifies research activities along two dimensions—codifiability and tacit knowledge requirement—to identify a delegation boundary that is cognitive, not sequential: it cuts through every stage of the research pipeline, not between stages. I argue that AI agents excel at speed, coverage, and methodological scaffolding but struggle with theoretical originality and tacit field knowledge. The paper concludes with an analysis of three implications for the profession—augmentation with fragile conditions, stratification risk, and a pedagogical crisis—and proposes five principles for responsible vibe researching.
@@ 1 Introduction
I address this question by developing a cognitive task framework and applying it to a concrete case: scholar-skill, a 23-skill plugin for Claude Code that covers the entire social science research workflow, from initial puzzle to journal submission. The system targets publication in top journals including the American Sociological Review, the American Journal of Sociology, Demography, Science Advances, Nature Human Behaviour, and Nature Computational Science.
@@ 2.2 From Chatbot to AI Agent
An AI agent, by contrast, is multi-step, persistent, and tool-using (Park et al., 2023; Wu et al., 2023). In the context of social science research, this means the agent can: read and write files to disk; execute R, Python, or Stata code via a shell; query databases (Zotero, NHANES, IPUMS); search the web and fetch papers; spawn parallel sub-agents for independent tasks; and invoke specialist skills with domain-specific knowledge bases. This is a qualitatively different architecture that enables a qualitatively different range of tasks (Shinn et al., 2023).
@@ 3.1 Architecture and Scope
Scholar-skill is a plugin for Claude Code—Anthropic’s command-line interface for the Claude language model—that organizes 23 specialist AI skills into a complete research pipeline. The system is installed locally, runs on the researcher’s laptop, and accesses local files, databases, and code environments. Table 2 summarizes the ten stage-groups and their constituent skills.
Table 2: Scholar-Skill: 23 Skills Organized by Research Stage
@@
Replication scholar-replication
Quality Assurance scholar-auto-improve
@@
Each skill is a self-contained workflow with its own knowledge base: reference documents, code templates, journal-specific norms, and quality checklists. The orchestrator—scholar-full-paper—coordinates all 23 skills across 17 phases with 43 quality gate items and three hard stops (at data safety, citation verification, and ethics compliance) that prevent the pipeline from advancing without meeting minimum standards.
@@ 3.2 Key Technical Capabilities
Idea formalization. Scholar-idea converts a broad research intuition into a formal, testable research question through a 10-step workflow: puzzle reframing, multi-angle exploration, variable mapping, hypothesis derivation, feasibility and novelty assessment, and a five-agent multi-agent evaluation panel (methodologist, theorist, domain expert, skeptic, and integrator) that stress-tests the proposed research question before the researcher commits to a design.
@@
Asset-driven writing. Scholar-write uses a three-tier knowledge graph built from 127 published papers: 32 by the author, 8 collaborator articles, and 87 top-journal exemplars. Tier 1 provides a section-snippets library organized by nine rhetorical functions. Tier 2 provides pre-extracted stylistic annotations (opening hooks, gap sentences, contribution claims, voice registers, citation density norms) for all 127 papers. Tier 3 provides full PDF access for deep reads. The system generates original prose that mimics the rhetorical structure of published papers while producing novel content—a distinction that separates structure mimicry from plagiarism. Each section undergoes a five-agent internal review panel—evaluating logic, rhetoric, journal fit, citation evidence, and accessibility—followed by a synthesizer that consolidates critiques with cross-agent agreement scoring and a reviser that implements changes, with user confirmation before advancing.
Peer review simulation. Scholar-respond spawns three or four reviewer agents in parallel—a methods specialist, a theory expert, a senior editor, and for computational papers, a fourth computational methods reviewer—each calibrated to the target journal’s priorities. The simulation produces a triage dashboard classifying concerns as MAJOR-FEASIBLE, MINOR-EASY, DISAGREE, or CONFLICT, with cross-reviewer overlaps flagged for priority attention.
@@
Replication package construction. Scholar-replication builds, documents, tests, and verifies a journal-ready replication package. It assembles scripts from the analysis pipeline into a structured directory with a master execution script, generates an AEA-template README with nine sections, runs clean-run validation in an isolated environment comparing reproduced outputs against originals, performs a paper-to-code correspondence audit mapping every table, figure, and in-text statistic to its producing script, and prepares archive deposits for Zenodo or Harvard Dataverse with DOI assignment. This distinguishes it from scholar-open, which produces open science declarations (templates, checklists); scholar-replication actively builds the package.
Continuous quality improvement. Scholar-auto-improve provides a self-improving quality engine operating in four modes: observe (monitor skill executions and log performance patterns), audit (systematically evaluate skill outputs against quality checklists), improve (propose and implement refinements to skill instructions, reference files, and quality gates), and evolve (identify gaps in the skill ecosystem and generate new capabilities). This meta-skill enables the system to learn from each research engagement, progressively tightening quality gates and expanding coverage—a capacity absent from static tool pipelines.
Multi-format manuscript output. The system produces manuscripts in four formats—Markdown, Word (.docx), , and PDF—via pandoc conversion, ensuring compatibility with diverse journal submission systems and co-author workflows.
@@ 6.2 The Stratification Risk
Cost. Claude Pro is twenty dollars a month. API tokens for a full 17-phase pipeline run cost five to fifteen dollars. For researchers at resource-poor institutions, these are meaningful barriers.
@@ 8 Discussion and Conclusion
First, AI agents are qualitatively different from prior tools. They execute multi-step reasoning workflows—not just computation or text—across the entire research pipeline. The scholar-skill system demonstrates that a single plugin can now cover 23 distinct research tasks from idea formalization to journal submission, coordinated by an orchestrator with quality gates and hard stops. The change in what a solo researcher can accomplish is real and significant.